Human ABO blood groups are controlled by a single gene with three alleles IA, IB, and i. The three alleles contribute to four blood group phenotypes. In addition to the ABO blood groups, there are two other loci that control blood phenotypes, the MN blood group locus and the Rh factor locus. The MN blood group has two codominant alleles, LM and LN. The Rh factor locus also has two alleles, Rh+ and Rh-, where Rh+ is dominant. The two phenotypes at the Rh factor locus are Rh+ and Rh-

Suppose in a population, the frequency of the ABO allele IA is 0.36, and the 1 allele frequency is 0.26. In the same population, the frequency of the LM allele is 0.47, and the frequency of the Rh+ is 0.62.

What is the frequency of the genotype IB i LM LN Rh+ Rh+ if the population is in Hardy-Weinberg equilibrium?
1
Expert's answer
2020-11-26T10:06:16-0500

To calculate the genotypic frequency of IBi LMLN Rh+ Rh+, we need the allele frequencies of IB, i, LM, LN, and Rh.

Given:

IA = 0.36

IB = 0.26

LM = 0.47

Rh = 0.62

Frequencies of alleles:

i = 1 - (0.36+0.26) = 0.38

LN = 1 - 0.47 = 0.53

Frequency of genotype IBi "= 2 \times 0.26 \times 0.38 = 0.1976"

Frequency of genotype LMLN "= 2 \times 0.47 \times 0.53 = 0.2491"

Frequency of genotype Rh+ Rh+ "= 2 \times 0.62 \times 0.62 = 0.7688"

Frequency of genotype IBi LMLN Rh+ Rh+ = product of the frequencies of the individual genotypes

"= 0.1976 \times 0.2491 \times 0.7688 = 0.0378"

Answer: 3.7 %